# Instigating Cooperation among LLM Agents Using Adaptive Information Modulation

### v2

@@ 1 Introduction
Interactions within social and sociotechnical systems are frequently characterized by a delicate balance of cooperation and competition, often leading to complex social dilemmas [1, 2, 3, 4, 5]. The challenge of governing these systems effectively hinges on the ability to foster and sustain prosocial behavior, thereby enhancing both system-level efficiency and fairness. The importance of such governance frameworks is underscored by their potential to increase overall social welfare, making the study of prosocial behavior crucial in a variety of contexts [6, 7, 8, 9].
@@
In this paper, we propose that the advent of advanced large language models (LLMs) offers a transformative opportunity to develop more robust governance mechanisms for sociotechnical systems. Recent studies have demonstrated that LLMs are capable of capturing nuanced strategic decision-making behaviors in classic games such as the Prisoner’s Dilemma, Snowdrift, and Stag Hunt. Importantly, these behaviors are influenced not only by the payoff structure of the game but also by the contextual framing, suggesting that LLM agents could serve as more accurate proxies for human behavior [20]. By leveraging the computational power of traditional agent-based models (ABMs) alongside LLMs, we can facilitate RL training without the need for overly simplistic assumptions, thus resulting in more efficient dynamic governance schemes. Our proposed framework is designed to address two key objectives: creating dynamic governance mechanisms for human interactions using LLM agents as proxies and enhancing prosocial behavior and alignment in hybrid environments with both LLM and human agents.
@@ 2 Related Work
Mechanisms to Promote Cooperation under Social Dilemma
A significant body of literature has examined mechanisms that influence cooperation in strategic interactions involving social dilemmas. Reputation and reciprocity are well-established factors, with interventions that establish and reinforce them consistently promoting cooperative behavior [21, 22]. Punishment and reward systems have also proven effective, deterring defection while incentivizing prosocial actions [23, 24]. Additionally, the structure of interaction networks plays a crucial role, either by identifying network characteristics that support cooperation [25, 26] or by examining how changes in the network structure over time can influence cooperation [27, 28].
@@ 3.3 Pro-social Promoting Agent (PPA)
The pro-social promoting agent (PPA) serves as the governing entity within our framework and is trained using reinforcement learning. Its role is to enhance network-level social welfare (the sum of all SLA scores) by dynamically adjusting the information levels provided to each agent.

### v3

@@ 1 Introduction
Interactions within social and sociotechnical systems are frequently characterized by a delicate balance of cooperation and competition, often leading to complex social dilemmas [1, 2, 3, 4, 5]. The challenge of governing these systems effectively hinges on the ability to foster and sustain prosocial behavior, thereby enhancing both system-level efficiency and fairness. The importance of such governance frameworks is underscored by their potential to increase overall social welfare, making the study of prosocial behavior crucial in a variety of contexts [6, 7, 8, 9, 10].
@@
In this paper, we propose that the advent of advanced large language models (LLMs) offers a transformative opportunity to develop more robust governance mechanisms for sociotechnical systems. Recent studies have demonstrated that LLMs are capable of capturing nuanced strategic decision-making behaviors in classic games such as the Prisoner’s Dilemma, Snowdrift, and Stag Hunt. Importantly, these behaviors are influenced not only by the payoff structure of the game but also by the contextual framing, suggesting that LLM agents could serve as more accurate proxies for human behavior [21]. By leveraging the computational power of traditional agent-based models (ABMs) [22] alongside LLMs, we can facilitate RL training without the need for overly simplistic assumptions, thus resulting in more efficient dynamic governance schemes. Our proposed framework is designed to address two key objectives: creating dynamic governance mechanisms for human interactions using LLM agents as proxies and enhancing prosocial behavior and alignment in hybrid environments with both LLM and human agents.
@@ 2 Related Work
Mechanisms to Promote Cooperation under Social Dilemma and the role of Information Modulariotn
A significant body of literature has examined mechanisms that influence cooperation in strategic interactions involving social dilemmas. Reputation and reciprocity are well-established factors, with interventions that establish and reinforce them consistently promoting cooperative behavior [23, 24]. Punishment and reward systems have also proven effective, deterring defection while incentivizing prosocial actions [25, 26]. Additionally, the structure of interaction networks plays a crucial role, either by identifying network characteristics that support cooperation [27, 28, 29, 30] or by examining how changes in the network structure over time can influence cooperation [31, 32].
@@ 3.3 Pro-social Promoting Agent (PPA)
The pro-social promoting agent (PPA) serves as the governing entity within our framework and is trained using reinforcement learning, essentially using a two-layer governance scheme similar to [60, 61]. Its role is to enhance network-level social welfare (the sum of all SLA scores) by dynamically adjusting the information levels provided to each agent.